OpenStudy (anonymous):

If f(x)= x+1^-1 and g(x)=x-2 , what is the domain of f(x) divided by g(x)

OpenStudy (carolinar7):

what do you think

OpenStudy (anonymous):

honestly I don't know, I want to say all values of x but im not sure :(

OpenStudy (zzr0ck3r):

Can you write out \(\dfrac{f(x)}{g(x)}\)

OpenStudy (anonymous):

no I don't know how sorry

OpenStudy (anonymous):

can you guys help me pleaseeee

OpenStudy (hybrik):

x+1^-1, So your saying the whole quantity by -1, or just "1"

OpenStudy (carolinar7):

\[\frac{ x+1 }{ x-2 }\]

OpenStudy (anonymous):

yeah the exponent is -1

OpenStudy (zzr0ck3r):

its the inverse \((x+1)^{-1}=\dfrac{1}{x+1}\)

OpenStudy (carolinar7):

no

OpenStudy (carolinar7):

The exponent is only on 1 not (x+1)

OpenStudy (zzr0ck3r):

So you have \(\dfrac{(x+1)^{-1}}{x-2}=\dfrac{1}{(x+1)(x-2)}\). Now what numbers force us to divide by \(0\)?

OpenStudy (carolinar7):

\[Did she mean (x+1)^{-1}\]

OpenStudy (hybrik):

The zeroes of (x+1)(x-2)

OpenStudy (zzr0ck3r):

yes. there is no reason to raise something to 1

OpenStudy (hybrik):

Also known as (x+1)=0, (x-2)=0, Solve for X on both of them @idgm_idontgetmath

OpenStudy (carolinar7):

x=-1,2

OpenStudy (zzr0ck3r):

please let them figure something out man. you are not helping if you give thm everything....

OpenStudy (hybrik):

But im trying to show him to solve zeroes??????

OpenStudy (zzr0ck3r):

Not you.

OpenStudy (anonymous):

@carolinar7

OpenStudy (anonymous):

@vera_ewing

OpenStudy (anonymous):

@juanpabloJR

OpenStudy (anonymous):

@campbell_st can you help?

OpenStudy (campbell_st):

ok... here is what I'll do, if you are happy that \[\frac{f(x)}{g(x)} = \frac{1}{(x +1)(x-2)}\] you need to solve x + 1 = 0 and x - 2 = 0 can you do that

OpenStudy (campbell_st):

so here is an image of f(x)/g(x)

OpenStudy (campbell_st):

the curve continues to the left.. to- infinity and to the right to infinity however the solutions to x + 1 = 0 and x -2 = 0 create vertical asymptotes... this is where the cruve doesn't exist but you can see that the curve exists between the asymptotes so there are 3 regions where the curve exists or can be drawn...

OpenStudy (campbell_st):

so the regions start with - infinity to the 1st asymptote between the asymptotes and the right asymptote to infinity... this will be the domain

OpenStudy (anonymous):

If im understanding right the solution is B then

OpenStudy (campbell_st):

that's correct

OpenStudy (anonymous):

awww thanks